### v1

# Spatially resolved H emission in B14-65666: compact starbursts, ionizing efficiency and gas kinematics in an advanced merger at the Epoch of Reionization

11email: cprieto@cab.inta-csic.es 22institutetext: Departamento de FĂsica de la Tierra y AstrofĂsica, Facultad de Ciencias FĂsicas, Universidad Complutense de Madrid, E-28040, Madrid, Spain 33institutetext: Space Telescope Science Institute (STScI), 3700 San Martin Drive, Baltimore, MD 21218, USA 44institutetext: Department of Astronomy, Stockholm University, Oscar Klein Centre, AlbaNova University Centre, 106 91 Stockholm, Sweden 55institutetext: Centro de AstrobiologĂa (CAB), CSIC-INTA, Camino Bajo del Castillo s/n, E-28692 Villanueva de la Cañada, Madrid, Spain 66institutetext: Max Planck Institut fĂŒr Astronomie, Königstuhl 17, D-69117, Heidelberg, Germany 77institutetext: Kapteyn Astronomical Institute, University of Groningen, P.O. Box 800, 9700AV Groningen, The Netherlands 88institutetext: Cosmic Dawn Center (DAWN), Copenhagen, Denmark 99institutetext: Physikalisches Institut der UniversitĂ€t zu Köln, ZĂŒlpicher Str. 77, 50937 Köln, Germany 1010institutetext: European Space Agency, Space Telescope Science Institute, Baltimore, Maryland, USA 1111institutetext: DTU-Space, Elektrovej, Building 328, 2800, Kgs. Lyngby, Denmark 1212institutetext: DARK, Niels Bohr Institute, University of Copenhagen, 2200 Copenhagen, Denmark 1313institutetext: Telespazio UK for the European Space Agency (ESA), ESAC, Camino Bajo del Castillo s/n, 28692 Villanueva de la Cañada, Madrid, Spain 1414institutetext: European Space Agency (ESA), European Space Astronomy Centre (ESAC), Camino Bajo del Castillo s/n, 28692 Villanueva de la Cañada, Madrid, Spain 1515institutetext: School of Physics & Astronomy, Space Park Leicester, University of Leicester, 92 Corporation Road, Leicester LE4 5SP, UK 1616institutetext: Steward Observatory, University of Arizona, 933 North Cherry Avenue, Tucson, AZ 85721, USA 1717institutetext: Leiden Observatory, Leiden University, PO Box 9513, 2300 RA Leiden, The Netherlands 1818institutetext: Division of Physics, Faculty of Pure and Applied Sciences, University of Tsukuba,Tsukuba, Ibaraki 305-8571, Japan 1919institutetext: Tomonaga Center for the History of the Universe (TCHoU), Faculty of Pure and Applied Sciences, University of Tsukuba, Tsukuba, Ibaraki 305-8571, Japan 2020institutetext: National Astronomical Observatory of Japan, 2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan 2121institutetext: Waseda Research Institute for Science and Engineering, Faculty of Science and Engineering, Waseda University, 3-4-1 Okubo, Shinjuku, Tokyo 169-8555, Japan 2222institutetext: ETH ZĂŒrich, Institute for Particle Physics and Astrophysics, Wolfgang-Pauli-Str. 27, 8093 ZĂŒrich, Switzerland
Spatially resolved H emission in B14-65666: compact starbursts, ionizing efficiency and gas kinematics in an advanced merger at the Epoch of Reionization
We present MIRI/JWST medium resolution spectroscopy (MRS) and imaging (MIRIM) of B14-65666, a source identified as a Lyman-break and interacting galaxy at a redshift of â=â7.15. We detect the H line emission in this system, revealing a spatially-resolved structure of the H emitting gas, which consists of two distinct galaxies, E and W, at a projected distance of 0.4 arcsec apart (i.e. 2.2 kpc). One of the galaxies (E) is very compact (upper limit for the effective radius of 63 pc) in the rest-frame ultraviolet light, while the other galaxy (W) is more extended (effective radius of 348 pc), showing a clumpy structure reminiscent of a tidal tail. The total H luminosity implies that the system is forming stars at a rate of 76â±â8 Mââyr-1 and 30â±â4 Mââyr-1 for E and W galaxies, respectively. The ionizing photon production efficiency, ), for galaxies E and W has values of 25.1â±â0.1 Hz erg-1 and 25.5â±â0.1 Hz erg-1, within the range measured in galaxies at similar redshifts. The high values derived for the H equivalent widths (832â±â100 and 536â±â78 ) and the distinct locations of the E and W galaxies in the ) equivalent width (H) plane, indicate that the system is dominated by a young (less than 10 Myr) stellar population. The overall spectral energy distribution suggests that in addition to a young stellar population, the two galaxies may have mature (more than 100 Myr) stellar population and very different dust attenuation with galaxy E showing a larger attenuation (AVâ=â1.5 mag), compared to the almost dust-free (AVâ=â0.1 mag) galaxy W. The derived Star Formation Rate (SFR) and stellar masses identify the two galaxies as going through a starburst phase characterized by a specific SFR (sSFR) of 40âââ50âGyr-1. Galaxy E has an extreme stellar mass surface density (6ââ104âMââpc-2), close to that of the nuclei of low- galaxies while galaxy W (103âMâ pc-2) is consistent with the surface densities measured in galaxies at these redshifts. The kinematics of the ionized gas traced by the H line show a velocity difference of 175ââ28 kmâs-1 between the two components of B14-65666, and a broader profile for galaxy W (312â44 kmâs-1) relative to galaxy E (243ââ41 kmâs-1). The detailed study of B14-65666 shows the complex stellar and Interstellar Medium structure in merging galaxy systems was already in place by the Epoch of Reionization. The general properties of B14-65666 agree with those predicted for massive merging systems at redshifts of 7 and above in the FIRSTLIGHT cosmological simulations. The in-depth study of systems like B14-65666 reveal how galaxy mergers in the early Universe drive intense star formation, shape the interstellar medium, and influence the buildup of stellar mass, just 700âââ800 Myr after the Big Bang.
Key Words.:
Galaxies: high-redshift â Galaxies: starburst â Galaxies: ISM â Galaxies: interactions â Galaxies: individual: B14-656661 Introduction
The James Webb Space Telescope (JWST, Gardner et al. 2023), is transforming our understanding of galaxy formation, particularly during the so-called Epoch of Reionization (EoR). JWST has photometrically detected potential galaxy candidates up to redshift 25 (PĂ©rez-GonzĂĄlez et al. 2025), and spectroscopically confirmed galaxies up to redshift 14.44 (Naidu et al. 2025). At these redshifts, the rest-frame optical emission lines and continuum are redshifted into the mid-infrared (5âââ25âm), the spectral range covered by the Mid-Infrared Instrument (MIRI, Rieke et al. 2015; Wright et al. 2015, 2023). At redshifts of 7 and above, this instrument allows us to explore the strong optical emission lines, such as H.
Mergers are expected to have a greater impact on galaxy evolution at higher redshifts than at lower redshifts, as the merger rate increases with redshift (Duan et al. 2025). The gravitational forces from mergers can cause significant disruptions in galaxy structures and movements, leading to features such as tidal tails and asymmetric stellar envelopes (e.g.Toomre & Toomre 1972; Mihos & Hernquist 1996; Bushouse et al. 2002; Conselice et al. 2003; Spilker et al. 2022). These disturbances can cause gas to flow from the outer regions into the center of galaxies, triggering the formation of new stars in central areas (PĂ©rez-GonzĂĄlez et al. 2024c). Simulations such as FirstLight (Ceverino et al. 2017) show that about 10% of galaxies at redshifts 6âââ10 exhibit clumpy structures, primarily formed through mergers. The central clumps are primarily composed of older stellar populations exceeding 50âMyr, while off-center clumps originate from gas debris along tidal tails during major mergers. These mergers have a short life, typically merging into the main galaxy within a few tens of million years (Nakazato et al. 2024). Observational studies also support the presence of mergers in the EoR, with recent works identifying clear merger signatures in high-redshift galaxies (e.g. Ălvarez-MĂĄrquez et al. 2023; Arribas et al. 2024; Harshan et al. 2024; Hu et al. 2025).
High- star-forming galaxies (SFGs) are characterized by optical lines with high (rest-frame) values of the equivalent widths (EW up to 3000â) in extreme emission line galaxies (Boyett et al. 2024). There is also a trend of increasing EW with redshift which indicates an evolution in the galaxy specific star formation rate (Rinaldi et al. 2023; Rinaldi et al. 2025) and young stellar populations (ÂĄâ10âMyr). Studies at low redshifts have shown that galaxies in close pairs exhibit an SFR increase compared to isolated galaxies, highlighting the impact of gravitational interactions on gas dynamics and star formation (Barton et al. 2000). This enhancement is also observed at higher redshifts (6.5âââ8.5), with an excess in SFR of 0.26â±â0.11 dex found in very close pairs, with no significant increase at larger separations (Duan et al. 2024; PuskĂĄs et al. 2025).
B14-65666 (also known as the galaxy âBig Three Dragonsâ) was discovered based on wide-field imaging data of the UltraVISTA survey of the Cosmological Evolution Survey (COSMOS) field in Y, J, H and Ks near-infrared bands (Bowler et al. 2012, Bowler et al. 2014). It was then spectroscopically detected in Ly (Furusawa et al. 2016). An image captured with the Hubble Space Telescope (HST) showed that the galaxy consists of two clearly separated components in the rest-frame ultraviolet (UV), thus suggesting that the object could be undergoing a merger of two galaxies (Bowler et al. 2017). Observations performed with Atacama Large Millimeter/submillimeter Array (ALMA) with Band 6 reported the detection of a â160âm dust continuum at the position of the components (Bowler et al. 2018). It was also detected [OâIII]â88âm, [CâII]â158âm, and underlying dust continuum emission at â90âm, in Bands 6 and 8 (Hashimoto et al. 2019). Furthermore, they showed that [OâIII]â88âm and [CâII]â158âm can be decomposed into two components spatially resolved and kinematically separated. These components are associated with the two UV emission peaks identified in the HST images. Additional observations with ALMA Band 7 did not detect the presence of the far-infrared [NâII]â122âm fine-structure emission line, nor did ALMA Band 3 observations detect other lines such as CO (6-5), CO (7-6), and [CâI] (2-1) (Sugahara et al. 2021; Hashimoto et al. 2023).
Recent JWST Near Infrared Camera (NIRCam) imaging has revealed the complex morphology of the two galaxy components: galaxy E consists of a compact core, surrounded by diffuse, extended, rest-frame optical emission, and galaxy W has an elongated morphology interpreted as tidal tails (Sugahara et al. 2025). Follow-up observations with NIRSpec IFU from 2.9âââ5.3âm directly map the strength of the rest-frame optical emission lines across the source (Jones et al. 2024). The profile of these lines revealed the presence of both narrow and broad line emission, with the latter indicating tidal interaction or outflows. This analysis showed that both galaxies have subsolar metallicities (0.24âZâ and 0.3âZâ, E and W, respectively) and a high rate of star formation. The observations do not show evidence for the presence of an active galactic nucleus (AGN) in B14-65666.
B14-65666 is one of the few galaxies identified as undergoing a major merger during the Epoch of Reionization. This system allows the investigation of the physical properties of an early major merger, including its stellar structure, as well as the different phases of the Interstellar Medium (ISM). Additionally, this system provides a reference for direct comparisons with simulations. In this study, we present the first spatially resolved spectroscopy of the H emission in the system B14-65666 at â=â7.15 based on new MIRI/JWST integral field spectroscopy. The paper presents several of the physical properties of B14-65666 based on the combination of the new MIRI data with ancillary NIRCam imaging and ALMA spectroscopy. This paper is organized as follows. In Sect. 2 we present the observations and data calibrations. In Sect. 3 we outline our analysis methods of spectroscopy and photometry. In Sect. 4 we present the results of the analysis. This section includes the subsections in which we study the Spectral Energy Distribution of the galaxy (4.1), the attenuation (4.2), the star formation rate and burstiness (4.3), the H EW and ionizing photon production efficiency (4.4), the kinematics (4.5), stellar and dynamical mass (4.6), the mass-size relation (4.7), the starburst nature (4.8), and finally we analyze the origin and evolution of the system according to simulations (4.9). We discuss and analyze the spectral and spatial distribution of the H line and the derived properties of the galaxies and we study the morphology and SFR of the galaxy and compare it with local galaxies. Our conclusions are presented in Sect. 5.
In this paper we assume a flat CDM cosmology with â=â0.310, and â=â67.7âkmâs-1âMpc-1 (Planck Collaboration et al., 2020). With these parameters, 1ââ is equal to 5.275 kpc at â=â7.15 and the luminosity distance is =â72276.9 Mpc. For these cosmological parameters and redshift, the age of the Universe corresponds to 740 Myr. We use vacuum emission line wavelengths throughout the paper.
2 Observations and data calibrations
2.1 MIRI
B14-65666 was observed with MIRI on the 30th of April of 2023 as a part of the MIRI European Consortium guaranteed time observations (program ID 1284, PI: L. Colina). The observations are composed of MIRI imaging (MIRIM, Bouchet et al. 2015; Dicken et al. 2024) and integral field spectroscopy using the Medium Resolution Spectrograph (MRS, Wells et al. 2015; Argyriou et al. 2023).
2.1.1 MIRI spectroscopy
The MRS observations were performed with the SHORT band for channels 1âââ4. We only use Channel 1 in this study. This channel covers the wavelengths 4.9âââ5.74âm, which targets the H emission line at redshift 7.15, and other fainter emission lines, such as [NâII]â6550,6585â and [SâII]â6718,6733â. The total on-source exposure time of the MRS observations is 22743 seconds. The positive and negative 4-point dither patterns for point sources are used. For each dither position, a total of 6 integrations are obtained using the SLOWR1 readout mode with 19 groups each. When the MRS is targeting B14-65666, a parallel field is imaged using MIRI F770W filter. This imaging data is utilized exclusively for the astrometric calibration of the MRS (see Sect. 2.4).
We use version 1.15.1 of the JWST calibration pipeline and context 1276 of the Calibration Reference Data System (CRDS) to process and calibrate the MRS observations of B14-65666. The standard MRS pipeline procedure (Bushouse et al. 2024) is used with additional customized steps to improve the quality of the final MRS calibrated products and background (see Ălvarez-MĂĄrquez et al. 2023, 2024 for details). Additionally, we implement the algorithm developed by Spilker et al. (2023) to correct the striping pattern produced by the cosmic ray showers in the spectral cube. The final Channel 1 cube has a spatial and spectral sampling of 0.13âââ0.13âââ0.8ânm (Law et al. 2023), and the resolving power () is about 3500 (Labiano et al. 2021; Jones et al. 2023).
2.1.2 MIRI imaging
The MIRIM observations of B14-65666 were performed using the F560W filter covering the rest-frame optical continuum at â0.7âm and the H emission line. The MIRIM F560W image has a total on-source integration time of 2675 seconds. The observational setup consists of a 4-point dither pattern, with the dither composed of 2 integrations of 120 groups using the FASTR1 readout mode.
The MIRIM F560W observations are calibrated with version 1.12.0 of the JWST pipeline (Bushouse et al., 2024) and context 1150 of the CRDS. We follow the general procedure to calibrate the data, but we also include more steps to correct for striping artifacts and background gradients (see PĂ©rez-GonzĂĄlez et al. 2024a; Ăstlin et al. 2025 for details). The final MIRIM F560W mosaic has a pixel size of 0.06ââ, after resampling, and a field of view (FoV) of 74âââ113ââ. The Point Spread Function (PSF) Full Width Half Maximum (FWHM) is 0.207ââ (Libralato et al. 2024).
2.2 NIRCam
The NIRCam (Rieke et al., 2023) data used in this work was presented in a previous study (Sugahara et al. 2025). The observations were taken with filters F115W, F150W, F200W, F277W, F356W and F444W (JWST GO1 program ID 1840; PIs: J. Ălvarez-MĂĄrquez and T. Hashimoto). At redshift 7.15, F115W, F150W and F200W sample the rest-frame ultraviolet spectrum, F277W and F356W cover the Balmer break region, and F444W targets the rest-frame continuum around 5000âĂ and the [OâIII]â4960,5008â emission lines.
NIRCam and MIRIM images present large variation on the spatial resolution, with PSF FWHMs increasing with wavelength from 0.04ââ to 0.21ââ. To obtain consistent multi-wavelength aperture photometry in the different NIRCam and MIRIM images, we perform a PSF homogenization in the NIRCam images to match the spatial resolution of NIRCam images to the coarser MIRIM F560W image. However, we also use the NIRCam images with the native spatial resolution to investigate the structure of B14-65666 in the rest-frame UV (see Sect. 3.4).
The PSF homogenization procedure includes the following steps. We employ the observed effective PSFs (ePSFs) for MIRI (Libralato et al. 2024) and for NIRCam (Nardiello et al. 2022)111https://www.stsci.edu/~jayander/JWST1PASS/LIB/PSFs/STDPSFs/. We compare these ePSFs with point-like sources from our observations, and conclude that they are consistent with each other. The first step is to rotate the NIRCam ePSF to the position angle (PA) of the MIRI image (PA=108Âș) ensuring orientation alignment between the two ePSFs. We rotate the MIRI ePSF to match the PA of the MIRI image. Next, we resample the pixel scale of the MIRI ePSF (0.028ââ, Libralato et al. 2024) to match that of the MIRI imager, and the NIRCam ePSF (0.008ââ for F115W, F150W, F200W and 0.016ââ for F277W, F356W, F444W, Nardiello et al. 2022) to match the pixel scale of the MIRI imager. Using these aligned and resampled ePSFs, we then generate a matching kernel using the python package PHOTUTILS to bridge the differences between the NIRCam and MIRI ePSFs. Finally, to apply the kernel using the python function creatematchingkernel (Bradley et al. 2022), we rescale it to the NIRCam image pixel size (0.03ââ after resampling). Subsequently, we convolve the NIRCam images with the kernel, thus homogenizing the PSF.
2.3 ALMA ancillary data
The ALMA data used in this work was previously presented in other studies. ALMA Bands 6 and 8 data were taken in Cycles 4 (ID 2016.1.00954.S, PI: A. K. Inoue) and 5 (ID 2017.1.00190.S, PI: A. K. Inoue) to target the [CâII]â158âm and [OâIII]â88âm emission lines (Hashimoto et al. 2019). ALMA Band 7 data was taken during Cycle 7 (ID 2019.1.01491.S, PI: A. K. Inoue; Sugahara et al. 2021). In this article, we use the spectra of [CâII]â158âm and [OâIII]â88âm, and the emission line and continuum (90, 120 and 160âm) fluxes presented in Hashimoto et al. (2019) and Sugahara et al. (2021).
2.4 MIRI-NIRCam astrometry
We correct the astrometry in the MIRI F560W image by using the centroid of two point-like sources within the FoV of the available GAIA DR3 (Gaia Collaboration et al., 2023) coordinates. The uncertainty in the final positioning is 18 mas. With the corrected MIRIM F560W image, we align the simultaneous MIRIM F770W image with two point-like sources in the common FoV of both images. As the MRS data and the MIRIM F770W image were taken simultaneously, we use the astrometric offset found in the MIRIM F770W image to correct the MRS 1SHORT cube, getting an uncertainty of less than 60âmas (0.3âkpc). Finally, we align the NIRCam images using the astrometrically corrected MIRIM F560W image as for the simultaneous MIRIM F770W image. The NIRCam images have an uncertainty smaller than 35âmas, obtained with the available star and galaxy that have GAIA astrometry (Gaia Collaboration et al., 2023).
3 Analysis
3.1 MIRI and NIRCam images and H line map
Previous HST images showed that B14-65666 is composed of two spatially separated galaxies in the rest-frame UV, a brighter galaxy in the northeast (galaxy E) and a fainter galaxy in the southwest (galaxy W) (Bowler et al. 2017). ALMA data further demonstrated that the [OâIII]â88âm and [CâII]â158âm emission lines can be spatially and spectrally decomposed into two components associated with galaxies E and W (Hashimoto et al. 2018). NIRCam images revealed the complex morphology of the individual galaxy components (Sugahara et al. 2025): galaxy E has a compact core surrounded by diffuse extended optical emission, while galaxy W displays elongated clumpy morphology.
MIRIM imaging and the MRS cover the emission line of H together with the continuum at 0.7âm. MIRI observations show a similar trend as NIRCam images. MRS shows an extended H emission associated with E and W galaxy components, additionally, each of these galaxy components is spectroscopically resolved as already presented in ALMA observations (Hashimoto et al. 2019). We present MIRI and NIRCam images in Fig. 1.
For this study, we define a total elliptical aperture for the whole system (named âgalaxies E+Wâ) with semi-major axis of 0.456ââ and semi-minor axis of 0.343ââ, and position angle (PA) of 85.83Âș. In the case of the galaxy E we define an elliptical aperture with semi-major axis of 0.211ââ and semi-minor axis of 0.201ââ, position angle (PA) of 112.83Âș. For the galaxy W we extract an ellipse with 0.228ââ and 0.177ââ semi-major and semi-minor axis respectively, PA of 112.83Âș. The coordinates of the center are shown in Table 2. The E and W elliptical apertures are compatible with the main emission of these galaxy components in our MIRIM F560W image. âGalaxies E+Wâ elliptical aperture covers the extended H emission from the MRS H line map. The H line map is generated by integrating the H line emission, centered at the H peak of galaxies E+W (5.3510âm) in the velocity range -500âÂĄâvâ[kmâs-1]âÂĄâ300. See Fig. 1.
Parameter | Galaxies E+W | Galaxy E | Galaxy W |
RA (ICRS) | 10:01:40.6858s | 10:01:40.6984s | 10:01:40.6707s |
DEC (ICRS) | +1:54:52.553s | +1:54:52.616s | +1:54:52.481s |
Redshift | |||
FWHM [kmâs-1] | |||
Flux H [ erg s-1 cm-2] | |||
H peak [m] | |||
Luminosity H [ erg s-1] | |||
Flux MIRIM F560W [Jy] | |||
Flux F115W [Jy] | |||
Flux F150W [Jy] | |||
Flux F200W [Jy] | |||
Flux F277W [Jy] | |||
Flux F356W [Jy] | |||
Flux F444W [Jy] |
3.2 H spectra and line fluxes
We extract the integrated H spectra of galaxies E+W, E and W using the three defined apertures (Fig. 1). We correct the two individual spectra for aperture losses, assuming that the two galaxies are unresolved sources for the MRS angular resolution taking into account the MRS PSF (Argyriou et al. 2023), which has a FWHM of 0.27ââââ0.35ââ at 5.35âm. The aperture correction is 1/0.57 both for galaxies E and W. In the case of the galaxies E+W, we also correct for aperture losses, assuming that the flux is mainly dominated by galaxies E and W, which are unresolved. The aperture correction in this case is 1/0.81. We also correct the spectrum for the contamination of flux from galaxy E in the W aperture by modeling galaxy E at the center of the aforementioned PSF and measuring how much of its flux would appear in the W aperture. We apply the same procedure for galaxy W to estimate its contribution to the E aperture. These corrections are 1/0.02 and 1/0.03, respectively.
We extract twelve 1D spectra using the same apertures as the galaxies E+W, E and W respectively at random positions of the MRS FoV clean of emission. We combine these spectra to generate the 1D median of the local background. The median is subtracted from the H spectra with the aim of removing any systematic residual feature left in the MRS calibration process. The three spectra are presented in Fig. 2.
The H spectrum is modeled by a single Gaussian function, to fit the emission line, and a second-order polynomial, to fit any residual background gradient. Using a Monte Carlo simulation, we estimate the uncertainties for the derived emission line parameters (such as FWHM, flux, and central wavelength). To determine the noise level in the spectrum, the root mean square (rms) is calculated from the continuum adjacent to the emission line. Using this rms value as the 1 standard deviation, we generate 1000 synthetic spectra by adding random Gaussian noise to the original spectrum. Each synthetic spectrum is then refitted to obtain new measurements of the line parameters. The final uncertainty for each parameter is given by the standard deviation of these measurements across all simulations. The Gaussian fit is shown in Fig. 2 and the results are presented in Table 2.
From the H line, the galaxies E+W of B14-65666 is at a redshift of 7.1513ââ0.0007. Galaxy E is redshifted to 7.1529ââ0.0004, however, galaxy W is blueshifted to 7.1481ââ0.0007. This implies that galaxy E and W present a velocity offset between the two components of 175ââ28 kmâs-1, compatible with that observed from [OâIII]â88âm and [CâII]â158âm, 177ââ16 kmâs-1 (Hashimoto et al. 2019).
The H line is spectrally resolved with the MIRI-MRS. It presents an intrinsic FWHM of 407ââ86 kmâs-1 for galaxies E+W and 243ââ41âkmâs-1 and 312ââ44âkmâs-1 respectively for E and W, after deconvolving with the Line Spread Function of the MRS (Labiano et al. 2021). Due to the offset in velocities the galaxies E+W component presents a higher FWHM than galaxy E and W. The observed H fluxes are for E+W, E and W, (23.1ââ3.1)â, (11.9ââ1.3)â, and (8.6ââ1.2)ââerg s-1âcm-2, respectively. All results are presented in Table 2. The H flux for galaxies E+W is in agreement within the uncertainties of the sum of the individual galaxy fluxes. This suggests that the H emission is mainly dominated by the two galaxies excluding any relevant diffuse component.
3.3 MIRIM and NIRCam photometry
MIRIM F560W photometry for galaxies E+W and for galaxies E and W is derived with the same apertures as for the H spectra (see Fig. 1 and Sect. 3.1). The local background level for each aperture is estimated in an elliptical annulus centered at the same position as the galaxies E+W elliptical aperture and with an inner radius of 0.95ââ and outer radius of 1.95ââ. Additionally, we use 300 apertures on blank regions of the MIRI image with the same size as the galaxy apertures to estimate the flux uncertainties. We also perform the NIRCam photometry using the same aperture as for MIRIM F560W. To do so, we utilize the NIRCam images (F115W, F150W, F200W, F277W, F356W, F444W) that are homogenized to the MIRI F560W PSF. The final observed MIRIM F560W and NIRCam fluxes are presented in Table 2. These NIRCam fluxes are lower by a factor of than the ones obtained from previous result (Sugahara et al. 2025) because of the PSF homogenization from NIRCam F444W to MIRI F560W (see Sect. 2.2) and the different aperture sizes.
Galaxy E is an unresolved source in the NIRCam F150W image with an estimated physical size of 63 pc (see Sect. 3.4 for details). The selected aperture mainly contains emission from galaxy E, thus, we implement aperture correction in the NIRCam and MIRIM photometry using the MIRIM F560W PSF. This aperture correction has a value of 1/0.67. On the contrary, galaxy W is fully resolved in the NIRCam F150W, and for that reason, we do not implement aperture correction to its photometry.
3.4 Stellar structure and effective radius
We study the stellar structure of the two galaxies, E and W, individually. To achieve this, we utilize the F150W image, as it offers the best combination of signal to noise ratio and spatial resolution. It presents the largest exposure time, 1675âs, arriving at a sensitivity of 7.7ânJy at 5â depth, which is a factor of 2âââ3 better than any other NIRCam band. Also, the PSF FWHM is 0.067ââ, allowing us to resolve structure up to 0.35âkpc at redshift 7.15 (Sugahara et al. 2025). F150W image traces the UV emission, dominated by the young and attenuated stellar population. We work with native-resolution images (before PSF-matching and resampling), and the analysis is done using the code GALFIT (Peng et al. 2002).
Galaxy E consists of a compact core, which appears unresolved, and a non-dominant extended component. To obtain the effective radius of galaxy E, we model the F150W data in GALFIT with a SĂ©rsic profile (with n=0.5, which gives a Gaussian profile) to fit the diffuse emission and a point-like component (using our ePSF) to fit the more compact part of the galaxy (Fig. 3). The surface brightness of galaxy E is compatible with a nuclear, point-like source plus a diffuse emission component at radii larger than 0.05ââ. However, as the core remains unresolved in the F150W image, we extend our analysis to the F115W image, in which the SNR is lower, but the spatial resolution is higher than in the F150W image. The core of the galaxy E also remains unresolved in the F115W image, thus we obtain an upper limit for the effective radius of 63âpc (0.012ââ).
Galaxy W has a clumpy elongated morphology and it is resolved in the F150W filter. It is well modeled by four different clumps. Three of them are unresolved point sources, which we model in GALFIT with point-like components. The other one is a more diffuse clump, associated with the brightest emission, modeled with a point-like and a SĂ©rsic profile (with an index nâ=â0.5) components (Fig. 3). When considering the galaxy W as a single entity and circularizing the profile, the intrinsic effective radius is 348âpc (0.057ââ), after deconvolving with the NIRCam F150W PSF.
The merger nature of B14-65666 is supported by the morphological analysis using the statmorph code (Rodriguez-Gomez et al. 2019). We analyze galaxies E and W separately and find that in both cases, the shape asymmetry, AS â a reliable tracer of merging galaxies (Rinaldi et al. 2024a) â is greater than 0.2, reinforcing the idea that each galaxy exhibits significant asymmetry. Specifically, we find AS of 0.35 and 0.80 for galaxies E and W, respectively. We also test the multimode-intensity-deviation () statistics, which are useful for detecting multi-component systems, such as the double-nucleus of a late-stage merger, highly disordered post-merger remnants, galaxies with bright star-forming clumps in rest-UV emission or an apparently single galaxy with an extended emission component(s) (Freeman et al. 2013). (multimode statistics) quantifies how a galaxyâs light is distributed into distinct regions, and (deviation statistic) measures the offset between the brightest region and the overall centroid. values approaching 1 are likely to indicate a double-nucleus, while values close to zero are interpreted as a single source; values close to zero indicate symmetric and ordered morphologies such as spheroids and disks. In the case of galaxy E, ââ0.0 and ââ0.1. This corroborates the interpretation that galaxy E is a single (compact spheroidal) source. In the case of galaxy W, ââ0.4 and â0.3, indicating that galaxy W is composed of at least two clumps, with one of them dominating the emission.
4 Results and discussion
We detect and spatially resolve the rest-frame optical continuum and H emission in the system B14-65666 at â=â7.15 with MIRI imaging and spectroscopy. The new MIRI data are combined with ancillary NIRCam imaging and ALMA spectroscopy to obtain the physical properties of galaxies E and W, and of the entire system: star formation rate, instantaneous SFR, ionizing photon production efficiency, H equivalent width, dynamical mass, stellar mass surface density, and SFR surface density.
Parameter | Galaxies E+W | Galaxy E | Galaxy W | |
Mâ [ Mâtotal ](a) | ||||
Mdyn [ Mâtotal ]â | - | |||
Ageyoung [Myr](a) | ||||
Agemain [Myr](a) | ||||
Metallicity [Zâ](a) | ||||
EW0(H) []â | - | |||
EW0(H) [](a) | ||||
[âLâ](a) | ||||
[ W](a) | ||||
[mag](a) | ||||
[mag] | - | |||
[Hz erg-1])â | - | |||
SFR [Mâyr-1](a) | ||||
SFRâ[Mââyr-1]â | - | |||
SFR â[Mââyr-1]â | - | |||
SFRâ[Mââyr-1] | 140 130 | 68 64 | 53 50 | |
SFRâ[Mââyr-1]â | 107 | 50 | 54 | |
SFRUVâ[Mââyr-1]â | 72 4 | 41 3 | 37 2 | |
SFRUV+TIRâ[Mââyr-1]â | 147 | 66 | 88 |
4.1 Spectral energy distribution and star formation history
We perform a SED fitting using CIGALE (Noll et al. 2009; Boquien et al. 2019) for galaxies E+W as well as for the two individual components. The SED covers the 1âm to 1.3âmm observed wavelength range, combining JWST photometric data (NIRCam and MIRIM) with ALMA continuum imaging (Hashimoto et al. 2019; Sugahara et al. 2021) along with our H and [OâIII]â88âm (Hashimoto et al. 2019) line fluxes. The redshift used in the analysis is the one derived spectroscopically from the H line.
We have made different assumptions to represent the star formation history of the system. The best results are determined from CIGALE fitting by minimizing the 2 value and ensuring the best agreement with observed photometry and line fluxes. These are obtained when considering two components: an instantaneous burst with ages ranging from 30 to 500 Myr, to model the mature stellar population, and a constant burst with ages lower than 10 Myr, to model the young stellar population. We use the Bruzual & Charlot (2003) stellar population models and the Chabrier (2003) Initial Mass Function (IMF). The nebular continuum and emission lines are modeled using electron densities from 100 to 1000 cm-3 and the ionization parameter (logU) ranging from -3.5 to -1.5. The metallicity is treated as a free parameter, ranging from 0.03âZâ to 0.7âZâ, as previous studies showed that it varies in the two galaxies E and W (Sugahara et al. 2025; Jones et al. 2024). We use the dust attenuation law from Calzetti et al. (2000), and the dust emission is modeled using the Draine et al. (2014) models.
The best-fit models for the galaxies E+W and for galaxies E and W are shown in Fig. 4. These models reveal that B14-65666 is dominated by a young stellar burst characterized with a total SFR of 127ââ66âMyr-1 and a mass for the young stellar component of (8ââ1)ââ108 Mâ. Galaxies E and W have SFRs of 84ââ49 and 50ââ12âMyr-1, respectively, and show the same young stellar component with a stellar mass of (4ââ1)ââ108âMâ. The SED fitting results indicate the presence of a mature stellar population in B14-65666 with a stellar mass of around 0.5ââ2ââ109âMâ and average ages about 200 Myr. SED fitting results are presented in Table 2. We conclude from the SED fitting that the B14-65666 system consists of two galaxies experiencing starbursts characterized by a similar age (see Sect. 4.9 for a more detailed discussion). Also, despite the compactness of galaxy E (see Sect. 4.7), the SED based on the NIRCam and MIRI imaging, and covering the rest-frame 0.16âââ0.78âm, does not correspond to that of the average Little Red Dots at redshifts â6âââ8 (PĂ©rez-GonzĂĄlez et al. 2024b; Kokorev et al. 2024; Rinaldi et al. 2024a).
4.2 Nebular and stellar attenuation
Dust attenuation plays a key role in shaping the observed properties of B14-65666, as revealed by MIRI, NIRCam, and ALMA observations. Combining the MRS H and the NIRSpec H fluxes444We implement aperture corrections (ACâ=â1/0.79) using the NIRSpec IFS PSF to match the fluxes presented in Jones et al. (2024) to the larger aperture of our MRS measurements., the H/H ratio gives observed values of 4.5â0.6 and 2.9ââ0.4 for galaxies E and W, respectively. Note that these values correspond to the narrow component of the line emission for both galaxies as the broad component present in the H line (Jones et al. 2024) is not detected in H with the present data (see Sect. 4.5). Assuming case B recombination, Teââ1.5âââK and neâ=ââcm-3, the theoretical ratio is 2.8. Using Cardelli et al. (1989), we get that the nebular visual extinction () is â=â1.5ââ0.4 mag for galaxy E, and it is â=â mag for galaxy W. Thus the nebular attenuation is higher for galaxy E, in agreement, within the uncertainties, with the values of â=â1.6ââ0.3 mag for galaxy E and â=â0.7â mag for W from Sugahara et al. (2025) (obtained with SED fitting and applying the conversion factor to nebular attenuation, fâ=â0.44; Calzetti 1997).
We also calculate the dust attenuation in the FUV, , with the IR-to-UV luminosity ratio (IRXâ=â/), as this is a robust tracer of . Taking the value of â=â0.7 for the total galaxy (Sugahara et al. 2025), we adopt expression (5) from Meurer et al. (1999) assuming the bolometric correction for the total light emitted by dust is 1, and for the total light emitted by stars, UV, is 1.68 (Meurer et al. 1999). We get a value of â=â1.50, compatible with that obtained with our SED fitting for galaxies E+W, â=â1.41â mag (Calzetti et al. 2000).
4.3 Star formation rate and burstiness
We derive the recent (i.e. less than 10 Myr) star formation rate directly from our H measurements, assuming solar (to facilitate comparison with other star formation tracers at other wavelengths) and sub-solar metallicity (0.29 Zâ, Jones et al. 2024 and see Table 2). We apply the corresponding SFR-to-H conversion factor for each metallicity, which is approximately 1.7 times lower in the subsolar case (Reddy et al. 2018). The final SFR values are corrected for attenuation using the H/H ratios (see previous section). We obtain SFRs of () Mââyr-1 and () Mââyr-1 for galaxies E and W and solar (subsolar) metallicities, respectively.
Additional estimates of the SFR can be obtained directly from far-infrared emission lines, or using a combination of multi-wavelength tracers. The [CâII]â158â emission line is a tracer of SFR in low-redshift star-forming and composite SF+AGN galaxies (De Looze et al. 2014), confirmed for high-redshift galaxies (Schaerer et al. 2020; Carniani et al. 2020). Adopting the ancillary [CâII]â158âm luminosities derived from ALMA (Hashimoto et al. 2019), values similar to those derived from our H measurements are obtained, within the uncertainties (see Table 2). The large uncertainty in the value of the SFR derived from the [CâII]â158âm emission line is due to the nature of this emission which could include extended or diffuse regions not directly linked to active star formation. Also, to trace less massive stars, we use the SFR(UV), for solar metallicities and considering a constant SFH over 100 Myr (Calzetti 2013).
Other SFR estimates involve the use of hybrid indicators such as the UV+TIR (Total Infrared) luminosity (Kennicutt & Evans 2012), or H+TIR (CatalĂĄn-Torrecilla et al. 2015; Calzetti 2013), that capture both obscured and unobscured star formation over longer periods of time (up to 100 Myr). Assuming a simple energy balance (Kennicutt et al. 2009), and considering the far-UV luminosity derived from our SED fitting (â=â(1.15ââ0.11)ââ), and the total infrared luminosity â=â(1.05ââ0.2)ââ (Hashimoto et al. 2019), we get SFR consistent with our extinction-corrected H estimates. The SFR derived from the H and TIR luminosities are also in good agreement with those obtained from UV+TIR (see Table 2). The SFRs derived in this analysis (70âââ140âMââyr-1) are lower than the SED-based SFR reported by Sugahara et al. (2025), SFRâ=â225ââMââyr-1, and the (narrow+broad) H-based SFR from Jones et al. (2024), SFRHÎČâ=â308ââ39âMââyr-1. This difference is attributed to variations in aperture sizes where the contribution of the extended emission (H) is significant, as well as other factors such as different estimates of the internal dust attenuation. Additionally, in the case of the H-based SFR, the presence of both narrow and broad components in the H line emission also contributes to the higher SFR value by a factor 2 (Jones et al. 2024).
A useful diagnostic for assessing the recent star formation in high- galaxies is the ratio of the SFR(H)-to-SFR(UV), referred as the burstiness parameter and which probes star formation over the last 10âMyr relative to the last 100âMyr. If the ratios are above 1, it indicates that the galaxy is in a bursty phase (Atek et al. 2024). We get a value of the ratio SFR(H)-to-SFR(UV) for solar metallicities of 1.8ââ0.1 and 0.8ââ0.1 for galaxies E and W, respectively (we would expect similar values for subsolar metallicities). When compared to the range of burstiness ratios reported by Atek et al. (2024) (0.7âÂĄâzâÂĄâ1.5), values above or close to the unity are typical in low-mass galaxies (MâÂĄâMâ), and below unity in high-mass galaxies (MâÂżâMâ). Thus, our values suggest that galaxy E shows signs of bursty activity, forming stars at nearly twice the average rate of the past 100 Myr, it is in a mild bursty phase, while galaxy W is forming stars at a rate closer to this past 100 Myr, i.e. experiencing a more continuous burst. This aligns with the results from Navarro-Carrera et al. (2024), who analyzed H emitters across zââ2âââ6.5 using JADES data, observing a correlation between enhanced SFR(H)/SFR(UV) and EW0(H), reinforcing the scenario where galaxies with above-unity SFR(H)/SFR(UV) experience a bursty phase in the last âMyr.
4.4 Equivalent width of H and ionizing photon production efficiency
To derive the rest-frame equivalent width of H (EW0(H)) for galaxies E+W and galaxies E and W, we make use of the F560W and H fluxes measured in our MIRIM and the MRS observations, respectively. First, we obtain rest-frame optical continuum emission by subtracting the contribution of the H line (Sect. 3.2) to the F560W flux (Sect. 3.3). The EW0(H) is then derived as the ratio of the H to the continuum flux under the H line. We obtain rest-frame equivalent widths 832ââ100âĂ and 536ââ78 Ă for galaxies E and W, respectively. The difference in the EW0(H) of the two galaxies is attributed to a different star formation history in the two galaxies with ionizing, young stars, and more mature, non-ionizing, stellar populations contributing in different proportions to the continuum emission. According to our SED analysis, galaxy E, with a larger EW0(H) value, has a slightly younger stellar population (6 Myr) relative to galaxy W (8 Myr).
The ionizing photon production efficiency, ), is the ratio of the ionizing to the non-ionizing UV flux. Assuming recombination, ) is empirically obtained as the ratio of the number of ionizing photons () derived from the H line (after extinction correction) and the luminosity at 1500 Ă . Assuming zero escape fraction, a dust attenuation as derived in Sect. 4.2, and a temperature of 1.5â 104âK (Ălvarez-MĂĄrquez et al. 2024) the integrated H luminosity provides a total number of ionizing photons () equal to âphâs-1 and âphâs-1 for galaxies E and W, respectively. Different approaches to dust correction can yield significantly different values of introducing uncertainty in its determination and affecting the interpretation of its physical implications (Matthee et al. 2017). Thus, correcting the UV and H luminosities, we get ionizing photon production efficiencies ) equal to and Hz erg-1 for galaxies E and W, respectively. This difference in the between the two galaxies of the system can be interpreted as partly due to the age difference in the young stellar populations in the two galaxies. Models indicate that in addition to the stellar continuum, there is a strong nebular contribution due to the two-photon continuum at 1500. This contribution can be close to 50% of the total for very young populations (1 Myr), decreasing with age and becoming almost negligible for 10-Myr-old populations (Katz et al. 2024). Thus the impact of this nebular continuum would be to increase the 1500â luminosity and consequently decrease the corresponding by factors 0.15 to 0.04 for ages 1 to 8 Myr. In addition, residual differential extinction in the two galaxies could also play an effect due to uncertainties in the derivation of the extinction in the ultraviolet.
We present the previous results, in the âââEW0(H) plane (see Fig. 5). Additional samples of galaxies at redshifts 4 to 9 are also presented for comparison. These include Lyman alpha emitters (LAE) at redshifts 3.8-6 (Maseda et al. 2020; Ning et al. 2023; Lam et al. 2019; Harikane et al. 2018), H and [OâIII] emitters at redshifts 6.1 to 8 (Sun et al. 2023; Rinaldi et al. 2023), and the mean of Lyman Break galaxies at redshift 8 (Stefanon et al. 2022). Individual galaxies with resolved integral field spectroscopy are also shown (RXCJ0600-2007 at â=â6.07, GimĂ©nez-Arteaga et al. 2024 and MACS1149-JD1 at â=â9.1, Ălvarez-MĂĄrquez et al. 2024). While galaxies E and W are forming stars at a rate much higher than the average galaxy at redshift 7, they have and EW0(H) in the range of high- galaxies. These values are consistent with the predicted values for young stellar populations. The Binary Population and Spectral Synthesis model (BPASS, Eldridge & Stanway 2020) demonstrates that the ) and the EW0(H) of a single-aged, instantaneous-starburst binary population increases as the age of the population decreases (see Fig. 5), with values compatible with those observed for young populations with ages from 1 to 20 Myr. The production of ionizing photons also depends on the presence of massive stars and therefore the initial mass function (Stanway et al. 2016). Values for the logarithm of the photon production efficiency above 25.5 Hzâerg-1 can only be obtained with massive bursts, âMâ due to the stochasticity of the formation of massive stars (Stanway & Eldridge 2023). The galaxies in the B14-65666 system and the high-z galaxies identified with JWST should have higher young stellar masses.
4.5 Kinematics of the ionized ISM
The kinematics of the ionized gas can be measured with the H emission line profile in our MRS data. Based on the integrated H line, we measure a redshift of â=â7.1513ââ0.0007 for the B14-65666 system, compatible with the values obtained with the FIR emission lines [OâIII]â88âm and [CâII]â158âm (â=â7.1521ââ0.0004, Hashimoto et al. 2019), and optical [OâIII] line (â=â7.1517ââ0.0001, Jones et al. 2024). The two galaxies of the B14-65666 system have a relative velocity of 175ââ28âkmâs-1 along our line of sight as derived from the centroid of the H line emission, showing it is a galaxy merger.
The kinematics of the ionized gas traced by H appears to show differences between the two galaxies. Specifically, we find that the FWHM for galaxy W is higher than for galaxy E, with values of 312ââ44âkmâs-1, and 243ââ41âkmâs-1, respectively. These values agree within the uncertainties from those measured with the far-infrared emission lines, [CâII]â158âm and [OâIII]â88âm. The [CâII]â158âm emission has similar kinematics in both galaxies (FWHM of 288ââ41âkmâs-1 for galaxy W and 276ââ34âkmâs-1 555The FWHM value for clump E has been recalculated and differs from that reported in Hashimoto et al. (2019), as confirmed through private communication with the authors. for galaxy E). For [OâIII]â88âm, galaxy E shows a broader line (322ââ41âkmâs-1) than galaxy W (267ââ42âkmâs-1). In Fig. 6 we present the comparison of these three lines for galaxies E+W and E and W. The observed differences in the two galaxies indicate the complex structure of the velocity field in the different phases of the interstellar medium. While the H line traces the overall kinematics of the ionized gas, the [CâII]â158âm line is closely associated with the neutral gas and the [OâIII]â88 m with the high excitation gas. Simulations also show that different tracers are sensitive to different kinematic properties (Kohandel et al. 2024). In addition, the continuum imaging (Sugahara et al. 2025) of both galaxies shows a different morphology, galaxy E as a very compact source while galaxy W shows a more extended structure, in particular in the rest-ultraviolet light, reminiscent of tidal tails. Moreover, low spectral resolution (Rââ1000) integral field spectroscopy with NIRSpec suggests the presence of a narrow and a broad kinematic component in both galaxies and in all optical emission lines tracing the ionized gas (Jones et al. 2024). These two components are potentially associated with the motions in the host galaxy (narrow), and the presence of high-velocity gas outflows (broad). The width of the narrow component is only 144ââ11âkmâs-1 for galaxy E and 227ââ6âkmâs-1 for galaxy W. These are significantly narrower than the values derived from our H line and far-infrared emission lines, obtained with a better spectral resolution but lower SNR. We test whether the observed H line profile could be compatible with the combination of the narrow and broad (653ââ10âkmâs-1 for galaxy E and 755ââ23âkmâs-1 for galaxy W) line components identified with NIRSpec. Our tests (see Fig. 7) indicate that the narrow+broad (N+B) component could be slightly compatible with our observed H line profiles. However, the simulated H N+B line profile appears to have a flux excess in the spectral range of the line (i.e. residuals mostly below zero) for galaxy E while the faint broad component in galaxy W is close in velocity to the peak emission in galaxy E. Since the two galaxies are separated by only 0.44 arcsec, this secondary emission in W could be due to a contamination from the brighter emission from E (Jones et al. 2024). Further higher spectral resolution data with NIRSpec, and higher SNR MRS spectra, would be required to confirm the presence of such a broad line emission in the two galaxies.
The consistency of the trends between our results and the ancillary data provides robust evidence of the dynamic differences between galaxies E and W. The broader profiles observed in galaxy W for both H and [CâII]â158âm may indicate enhanced turbulence or kinematic effects, potentially driven by interactions or feedback processes, as well as a possibility of a connection with outflows. On the other hand, the higher FWHM in [OâIII]â88 m for galaxy E suggests that highly ionized gas dynamics differ from those of lower-ionization states, pointing to localized physical conditions or energetic phenomena. Also, dust attenuation could impact the interpretation, in particular in galaxy E where a large visual extinction is measured. The optical lines are more susceptible to dust extinction, which can affect measurements of line widths and kinematic features, whereas far-infrared lines are less impacted by such effects.
4.6 Stellar and dynamical mass
Although the line profiles of the different emission lines suggest that the gas is not fully (dynamically) relaxed, here we obtain an upper limit of the dynamical mass of each of the galaxies assuming the system is close to relaxation and characterized by the random motions traced by the ionized gas velocity dispersion, i.e. not dominated by rotational motions. Under this working hypothesis, we follow previous work in similar systems like the low-z luminous infrared galaxies (e.g. equation (1) in Bellocchi et al. 2013). The dynamical masses we derive are (1.5ââ0.3)âââ and (8.5ââ2.4)âââ for galaxies E and W, respectively assuming velocity dispersions of 103ââ17 kmâs-1 and 132ââ19âkmâs-1 (obtained with the FWHM from our H line), and the effective radii derived from our NIRCam images (see Sect. 3.4 for details). The total stellar masses are derived from the multi-wavelength SED fitting and correspond to (15ââ9)âââMâ and (8ââ4)ââ108âMâ for galaxies E and W. Thus, while there is a good agreement between the stellar and dynamical mass for galaxy E, an order of magnitude difference appears for galaxy W. The difference in galaxy W could be due to the presence of a dominant molecular gas component in this galaxy amounting to close to 90% of the total mass. However, the stellar structure shows the presence of extended emission consisting of star-forming clumps along tidal tails (Sugahara et al. 2025). This suggests that this galaxy is far from relaxed and therefore the estimated dynamical mass is highly uncertain.
4.7 Mass-size relation
The two galaxies of the B14-65666 system occupy a very different location in the stellar mass-size plane (see Fig. 8). Galaxy W with an effective radius of 348 pc and a stellar mass of (8ââ4) 108 Mâ, is consistent with the expected values according to the relations derived for high redshift galaxies from the JADES program (redshifts 4 to 10; Langeroodi & Hjorth 2023), and from the CEERS, PRIMER-UDS, and PRIMER-COSMOS surveys (redshifts 3 to 9; Allen et al. 2025). Therefore, galaxy W appears as an average galaxy at redshift of about 7. On the other hand, the second component of the system, galaxy E, appears as deviating from the standard relations, showing a very small, compact size (upper limit of 63 pc to the effective radius) for a stellar mass of 15 (â9)ââ108âMâ. This translates into a high stellar mass surface density (6âââMââpc-2 for E, versus 1â Mââpc-2 for W), close to the limit of the highest values (âMââpc-2) measured in the stellar clusters and nuclei of low-z galaxies (GrudiÄ et al. 2019). Galaxies as compact as galaxy E and with similar stellar surface densities are not uncommon at high- (Ălvarez-MĂĄrquez et al. 2025). Star-forming clumps and galaxies with effective radius in the 30 to 200 pc range and a wide range of mass (4ââ107 to 2ââ1010 Mâ) are already known at redshifts above 6 (see Fig. 8; Claeyssens et al. 2023; Baggen et al. 2023).
Thus, B14-65666 is identified as a system that consists of a pair of interacting galaxies, with similar stellar masses (within a factor of 2) but with one of the galaxies (E) being extremely compact and already showing stellar mass surface densities close to those measured in the nuclei of galaxies at low redshift. Even more, the galaxies of the B14-65666 system are much less massive (ââ20âââ30), and far more compact (ââ10âââ60) than the main-sequence and the most luminous star-forming galaxies at low redshift, with effective radius â3.5 kpc (Leroy et al. 2021), and â2.2 kpc (Bellocchi et al. 2022) respectively (see Fig. 8). Thus, the galaxies E and W appear as building up their stellar mass as they go through the advanced phases of the interaction/merging process.
4.8 Starburst nature
The analysis of the SFR derived with H versus stellar mass plane reveals that both galaxies E and W are well above the main sequence (MS) of galaxies at redshift 7 (Heintz et al. 2023). Galaxies E and W of our system are situated between the region defined by Caputi et al. (2017) for H emitters, the starburst region (SB), and the MS (see Fig. 9). Such extremely compact starburst galaxies are expected to have short depletion times (¥100 Myr, see Franco et al. 2020) and therefore galaxy E may quench soon out of gas exhaustion - at least temporarily - which is in line with a bursty/stochastic SFH.
We compare the position in the SFR-stellar mass plot with more galaxies at similar redshifts (Fig. 9). The sSFR of E and W (51 and 38 Gyr-1 respectively) is similar to that of low-mass galaxies, but also to some very massive ones, such as REBELS-25 (sSFR=25 Gyr-1), a relaxed dynamically cold disc galaxy at similar redshift, â=â7.31 (Rowland et al. 2024). A recent large compilation of NIRSpec prism data and stacked spectra has shown that the average galaxy becomes progressively more bursty with increasing redshift as well as with increasing sSFR (Langeroodi & Hjorth 2024). Based on this, B14-65666 seems slightly less bursty (but consistent within error bars) than the average galaxy at similar redshift and sSFR. UV-faint galaxies at redshifts from 6 to 8 from the Cosmic Evolution Early Release Science (CEERS) JWST/NIRCam survey also show a bursty star formation history (Endsley et al. 2023). Galaxies from the ALMA-ALPINE [CâII] survey at redshifts 4 to 6 (Le FĂšvre et al. 2020; Faisst et al. 2020) show normal star formation following the main sequence of galaxies at these redshifts (Speagle et al. 2014). LIRGs at low redshift, with typical stellar masses in the range âMâ and SFR 30âMââyr-1 (Bellocchi et al. 2022) are placed above the low- main sequence plane (Elbaz et al. 2011) and indicating that at high redshifts LIRGs are less massive but more compact.
4.9 B14-65666: a merger of starburst galaxies in the early Universe
The picture that emerges from our analysis is that B14-65666 is a system of two galaxies in the process of interacting and merging. We have identified one of the components (E) as an extremely compact source, characterized by a stellar mass surface density close to that of nuclei of low-z galaxies. On the other hand, the second component of the system (galaxy W) shows a more elongated structure with four fainter clumps (Sugahara et al. 2025), reminiscent of star-forming regions along the tidal tails produced in gas-rich interacting galaxies. A significant velocity offset of 175âkmâs-1 between the two galaxies supports the merger scenario. Furthermore, the estimated stellar masses for each component are on the order of ââMâ, three to four orders of magnitude higher than typical stellar masses of star-forming clumps in galactic disks (ââââMâ, Adamo et al. 2024; Mowla et al. 2024), reinforcing the interpretation that we are observing two galaxy-scale systems. The stellar mass ratio between the components is â2:1, potentially consistent with a major merger. Additionally, we find differences in metallicity between the two components, with values derived with the SED-fitting of 0.24âZâ for E and 0.31âZâ for W, in agreement with line diagnostics from NIRSpec IFS (Jones et al. 2024). Collectively, the presence of tidal features, velocity offset, distinct stellar populations and metallicities, and the high stellar masses all provide compelling evidence that B14-65666 is an example of a merging system of two galaxies. In addition, galaxies E and W are experiencing a bursty phase with a SFR well above that of the main-sequence star-forming galaxies of the same mass at those redshifts.
Recent simulations of the formation of galaxies during the Epoch of Reionization (Nakazato et al. 2024 and references therein) show that systems identified as clumpy, i.e. with two or more high surface density star-forming clumps as traced by emission lines, represent 10 of all simulated sources. Moreover, most of the clumpy systems are produced by mergers, with the young stellar populations forming along the gas-rich tidal tails, and with specific SFR of the order of 50 Gyr-1. However, these clumpy systems appear to be temporal and short-lived, merging into a single galaxy in about 50 Myr (Hashimoto et al. 2023). Under this scenario, the compact, and highly dense, galaxy E could already be a protobulge produced by the merger of previous clumps while galaxy W would be in the process of merging of smaller, less massive, clumps before the two main components, E and W, merge together in a single, more massive galaxy at a later time, i.e. lower redshifts (see Figure 5 in Nakazato et al. 2024).
B14-65666 is not a unique system at those redshifts. Recent JWST observations have discovered merging systems of star-forming galaxies such as the interacting system SPT0311-58, which resides in the core of an extremely massive protocluster at â=â6.9 (Arribas et al. 2024), the source A2744-z7p9OD, the most distant protocluster at â=â7.88 (Hashimoto et al. 2023), or the JWSTâs Quintet, a merger of at least five galaxies containing over 17 clumps, at â=â6.7 (Hu et al. 2025). Detection and further studies with JWST of new systems like these will allow us to understand the formation process of massive galaxies in the early Universe.
5 Conclusions and summary
We present mid-infrared imaging and spectroscopy of the â=â7.15 Lyman-break galaxy B14-65666 taken with MIRI/JWST. The data provide the first detection of the H emission line of this source, a system of two merging galaxies at the Epoch of Reionization. The study also combines the new MIRI data with NIRCam data (Sugahara et al. 2025) and archival high angular resolution ALMA imaging (Hashimoto et al. 2018) to derive additional physical properties. The main results are summarized below:
âą
The H emitting gas shows a structure dominated by two spatially resolved sources (E and W) separated by a projected distance of 0.4 arcsec (i.e. 2.2âkpc) and relative velocity of 175âkmâs-1. The rest-frame ultraviolet light shows galaxy E as a very compact, unresolved source (upper limit for the effective radius of 63âpc) while galaxy W is more extended (348âpc) with the presence of fainter sources along an elongated structure.
âą
The SFR derived from the H luminosity is 76 and 30 Mââyr-1 (for solar metallicity as reference), for galaxies E and W respectively. Other measurements of the SFR using hybrid tracers, H and the luminosity TIR, and the UV plus TIR luminosities give values of 50 and 54 Mââyr-1, and 66 and 88 Mââyr-1 for galaxies E and W, respectively. These values are compatible within the uncertainties with the SFR derived from H. The ratio of H-to-UV based SFR (burstiness) indicates that while galaxy E is in a mild bursty phase (ratio of 1.8), galaxy W is experiencing a more continuous burst (ratio of 0.8).
âą
The equivalent width of the H line (832â±â100 and 536â±â78 , for galaxies E and W, respectively) and the ionizing photon production efficiency (â=â25.1â±â0.1 Hzâerg-1 and â=â25.5â±â0.1 Hzâerg-1) in the two galaxies are within the range of values measured in other samples of galaxies at these redshifts. Their different locations in the production efficiency âââequivalent width plane, with galaxy W above and E below the relation for high- galaxies is interpreted as due to slightly different ages in their young stellar populations.
âą
The specific star formation rate (40âââ50âGyr-1) indicates that the two galaxies involved in the B14-65666 system are experiencing an intense starburst phase, well above the typical values for galaxies of similar mass at those redshifts, and similar to less massive (â10) galaxies.
âą
The stellar mass surface density of galaxy E is extreme (â6âââMââpc-2) and close to the value measured in the nuclei of low- galaxies while galaxy W has a surface density (âMââpc-2) similar to that of high- galaxies.
âą
The H emission line profile shows a different gas kinematics, with galaxy W showing larger velocities (i.e. broader line) than galaxy E. The ionized gas kinematics traced by the H line shows differences with respect to the neutral and highly ionized gas traced by the far-infrared [CâII]â158âm and [OâIII]â88âm lines, respectively. This indicates a complex structure in the interstellar medium, likely due to the internal substructure of the young/mature stellar populations and dust distribution in these galaxies on (sub)kpc scales.
In summary, B14-65666 is a system consisting of two galaxies that are experiencing an extreme starburst phase due to the interaction/merging process. One of the galaxies (E) is already very compact and massive, with a stellar density close to that of the mature nucleus/bulge of low galaxies, while the other (W) is consistent with a more extended star formation, distributed in a few fainter clumps along a tidal tail. These results agree with recent cosmological simulations of clumpy merging systems in the early Universe (Nakazato et al. 2024). According to these simulations, this clumpy phase identified in B14-65666 is short-lived, with the bursty clumps merging in a short time scale (50 Myr) and forming a massive, single-nucleus galaxy at later times. Detailed studies of systems similar to B14-65666 will help understanding the physical conditions and mechanisms involved in the formation process of massive galaxies and their evolution in the early Universe, just a few hundred Myr after the Big Bang, and during the Epoch of Reionization.
Acknowledgements.
We thank Michele Perna and Gareth Jones for discussions on B14-65666. CPJ, JAM, LC acknowledge support by grant PIB2021-127718NB-100 from the Spanish Ministry of Science and Innovation/State Agency of Research MCIN/AEI/10.13039/501100011033 and by âERDF A way of making Europeâ. ACG acknowledges support by JWST contract B0215/JWST-GO-02926. AB, JM and GĂ acknowledge support from the Swedish National Space Administration (SNSA). KIC acknowledges funding from the Dutch Research Council (NWO) through the award of the Vici Grant VI.C.212.036 and funding from the Netherlands Research School for Astronomy (NOVA). AAH acknowledges support from from grant PID2021-124665NB-I00 funded by the Spanish Ministry of Science and Innovation and the State Agency of Research MCIN/AEI/10.13039/501100011033 and ERDF A way of making Europe. This work was supported by research grants (VIL16599,VIL54489) from VILLUM FONDEN. JPP and TVT acknowledge financial support from the UK Science and Technology Facilities Council, and the UK Space Agency. The project that gave rise to these results received the support of a fellowship from the âla Caixaâ Foundation (ID 100010434). The fellowship code is LCF/BQ/PR24/12050015. PGPG and LC acknowledge support from grant PID2022-139567NB-I00 funded by Spanish Ministerio de Ciencia e InnovaciĂłn MCIN/AEI/10.13039/501100011033, FEDER Una manera de hacer Europa. EI acknowledges funding from the Netherlands Research School for Astronomy (NOVA). IJ acknowledge support from the Carlsberg Foundation (grant no CF20-0534) and the Cosmic Dawn Center is funded by the Danish National Research Foundation under grant No. 140. SG acknowledges financial support from the Villum Young Investigator grant 37440 and 13160 and the Cosmic Dawn Center (DAWN), funded by the Danish National Research Foundation under grant DNRF140. AE and FP acknowledge support through the German Space Agency DLR 50OS1501 and DLR 50OS2001 from 2015 to 2023. The work presented is the effort of the entire MIRI team and the enthusiasm within the MIRI partnership is a significant factor in its success. MIRI draws on the scientific and technical expertise of the following organisations: Ames Research Center, USA; Airbus Defence and Space, UK; CEA-Irfu, Saclay, France; Centre Spatial de LiĂ©ge, Belgium; Consejo Superior de Investigaciones CientĂficas, Spain; Carl Zeiss Optronics, Germany; Chalmers University of Technology, Sweden; Danish Space Research Institute, Denmark; Dublin Institute for Advanced Studies, Ireland; European Space Agency, Netherlands; ETCA, Belgium; ETH Zurich, Switzerland; Goddard Space Flight Center, USA; Institute dâAstrophysique Spatiale, France; Instituto Nacional de TĂ©cnica Aeroespacial, Spain; Institute for Astronomy, Edinburgh, UK; Jet Propulsion Laboratory, USA; Laboratoire dâAstrophysique de Marseille (LAM), France; Leiden University, Netherlands; Lockheed Advanced Technology Center (USA); NOVA Opt-IR group at Dwingeloo, Netherlands; Northrop Grumman, USA; Max-Planck Institut fĂŒr Astronomie (MPIA), Heidelberg, Germany; Laboratoire dâEtudes Spatiales et dâInstrumentation en Astrophysique (LESIA), France; Paul Scherrer Institut, Switzerland; Raytheon Vision Systems, USA; RUAG Aerospace, Switzerland; Rutherford Appleton Laboratory (RAL Space), UK; Space Telescope Science Institute, USA; Toegepast- Natuurwetenschappelijk Onderzoek (TNO-TPD), Netherlands; UK Astronomy Technology Centre, UK; University College London, UK; University of Amsterdam, Netherlands; University of Arizona, USA; University of Cardiff, UK; University of Cologne, Germany; University of Ghent; University of Groningen, Netherlands; University of Leicester, UK; University of Leuven, Belgium; University of Stockholm, Sweden; Utah State University, USA. A portion of this work was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration. We would like to thank the following National and International Funding Agencies for their support of the MIRI development: NASA; ESA; Belgian Science Policy Office; Centre Nationale DâEtudes Spatiales (CNES); Danish National Space Centre; Deutsches Zentrum fur Luft-und Raumfahrt (DLR); Enterprise Ireland; Ministerio De EconomĂa y Competitividad; Netherlands Research School for Astronomy (NOVA); Netherlands Organisation for Scientific Research (NWO); Science and Technology Facilities Council; Swiss Space Office; Swedish National Space Board; UK Space Agency. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to the Author Accepted Manuscript version arising from this submission. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy (AURA), Inc., under NASA contract NAS 5-03127 for JWST; and from the European JWST archive (eJWST) operated by the ESDC. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This research made use of Photutils, an Astropy package for detection and photometry of astronomical sources (Bradley et al., 2022).References
- Adamo et al. (2024) Adamo, A., Bradley, L. D., Vanzella, E., et al. 2024, Nature, 632, 513
- Allen et al. (2025) Allen, N., Oesch, P. A., Toft, S., et al. 2025, A&A, 698, A30
- Ălvarez-MĂĄrquez et al. (2024) Ălvarez-MĂĄrquez, J., Colina, L., Crespo GĂłmez, A., et al. 2024, A&A, 686, A85
- Ălvarez-MĂĄrquez et al. (2025) Ălvarez-MĂĄrquez, J., Crespo GĂłmez, A., Colina, L., et al. 2025, A&A, 695, A250
- Ălvarez-MĂĄrquez et al. (2023) Ălvarez-MĂĄrquez, J., Crespo GĂłmez, A., Colina, L., et al. 2023, A&A, 671, A105
- Argyriou et al. (2023) Argyriou, I., Glasse, A., Law, D. R., et al. 2023, A&A, 675, A111
- Arribas et al. (2024) Arribas, S., Perna, M., RodrĂguez Del Pino, B., et al. 2024, A&A, 688, A146
- Atek et al. (2024) Atek, H., Labbé, I., Furtak, L. J., et al. 2024, Nature, 626, 975
- Baggen et al. (2023) Baggen, J. F. W., van Dokkum, P., Labbé, I., et al. 2023, ApJ, 955, L12
- Barton et al. (2000) Barton, E. J., Geller, M. J., & Kenyon, S. J. 2000, ApJ, 530, 660
- Bellocchi et al. (2013) Bellocchi, E., Arribas, S., Colina, L., & Miralles-Caballero, D. 2013, A&A, 557, A59
- Bellocchi et al. (2022) Bellocchi, E., Pereira-Santaella, M., Colina, L., et al. 2022, A&A, 664, A60
- Boquien et al. (2019) Boquien, M., Burgarella, D., Roehlly, Y., et al. 2019, Astronomy & Astrophysics, 622, A103
- Bouchet et al. (2015) Bouchet, P., GarcĂa-MarĂn, M., Lagage, P. O., et al. 2015, PASP, 127, 612
- Bowler et al. (2014) Bowler, A., Dunlop, J., McLure, R., et al. 2014, Monthly Notices of the Royal Astronomical Society, 440, 2810
- Bowler et al. (2018) Bowler, R. A. A., Bourne, N., Dunlop, J. S., McLure, R. J., & McLeod, D. J. 2018, Monthly Notices of the Royal Astronomical Society, 481, 1631
- Bowler et al. (2012) Bowler, R. A. A., Dunlop, J. S., McLure, R. J., et al. 2012, Monthly Notices of the Royal Astronomical Society, 426, 2772
- Bowler et al. (2017) Bowler, R. A. A., Dunlop, J. S., McLure, R. J., & McLeod, D. J. 2017, MNRAS, 466, 3612
- Boyett et al. (2024) Boyett, K., Bunker, A. J., Curtis-Lake, E., et al. 2024, MNRAS, 535, 1796
- Bradley et al. (2022) Bradley, L., SipĆcz, B., Robitaille, T., et al. 2022, astropy/photutils: 1.5.0
- Bruzual & Charlot (2003) Bruzual, G. & Charlot, S. 2003, MNRAS, 344, 1000
- Bushouse et al. (2024) Bushouse, H., Eisenhamer, J., Dencheva, N., et al. 2024, JWST Calibration Pipeline
- Bushouse et al. (2002) Bushouse, H. A., Borne, K. D., Colina, L., et al. 2002, The Astrophysical Journal Supplement Series, 138, 1
- Calzetti (1997) Calzetti, D. 1997, in American Institute of Physics Conference Series, Vol. 408, The ultraviolet universe at low and High redshift, ed. W. H. Waller (AIP), 403â412
- Calzetti (2013) Calzetti, D. 2013, in Secular Evolution of Galaxies, ed. J. Falcón-Barroso & J. H. Knapen, 419
- Calzetti et al. (2000) Calzetti, D., Armus, L., Bohlin, R. C., et al. 2000, ApJ, 533, 682
- Caputi et al. (2017) Caputi, K. I., Deshmukh, S., Ashby, M. L. N., et al. 2017, ApJ, 849, 45
- Cardelli et al. (1989) Cardelli, J. A., Clayton, G. C., & Mathis, J. S. 1989, ApJ, 345, 245
- Carniani et al. (2020) Carniani, S., Ferrara, A., Maiolino, R., et al. 2020, MNRAS, 499, 5136
- Catalån-Torrecilla et al. (2015) Catalån-Torrecilla, C., Gil de Paz, A., Castillo-Morales, A., et al. 2015, A&A, 584, A87
- Ceverino et al. (2017) Ceverino, D., Glover, S. C. O., & Klessen, R. S. 2017, MNRAS, 470, 2791
- Chabrier (2003) Chabrier, G. 2003, PASP, 115, 763
- Claeyssens et al. (2023) Claeyssens, A., Adamo, A., Richard, J., et al. 2023, MNRAS, 520, 2180
- Conselice et al. (2003) Conselice, C. J., Bershady, M. A., Dickinson, M., & Papovich, C. 2003, AJ, 126, 1183
- De Looze et al. (2014) De Looze, I., Cormier, D., Lebouteiller, V., et al. 2014, A&A, 568, A62
- Dicken et al. (2024) Dicken, D., MarĂn, M. G., Shivaei, I., et al. 2024, A&A, 689, A5
- Draine et al. (2014) Draine, B. T., Aniano, G., Krause, O., et al. 2014, ApJ, 780, 172
- Duan et al. (2025) Duan, Q., Conselice, C. J., Li, Q., et al. 2025, MNRAS
- Duan et al. (2024) Duan, Q., Li, Q., Conselice, C. J., et al. 2024, arXiv e-prints, arXiv:2411.04944
- Elbaz et al. (2011) Elbaz, D., Dickinson, M., Hwang, H. S., et al. 2011, A&A, 533, A119
- Eldridge & Stanway (2020) Eldridge, J. J. & Stanway, E. R. 2020, arXiv e-prints, arXiv:2005.11883
- Endsley et al. (2023) Endsley, R., Stark, D. P., Whitler, L., et al. 2023, MNRAS, 524, 2312
- Faisst et al. (2020) Faisst, A. L., Schaerer, D., Lemaux, B. C., et al. 2020, ApJS, 247, 61
- Franco et al. (2020) Franco, M., Elbaz, D., Zhou, L., et al. 2020, A&A, 643, A30
- Freeman et al. (2013) Freeman, P. E., Izbicki, R., Lee, A. B., et al. 2013, Monthly Notices of the Royal Astronomical Society, 434, 282
- Furusawa et al. (2016) Furusawa, H., Kashikawa, N., Kobayashi, M. A. R., et al. 2016, The Astrophysical Journal, 822, 46
- Gaia Collaboration et al. (2023) Gaia Collaboration, Vallenari, A., Brown, A. G. A., et al. 2023, A&A, 674, A1
- Gardner et al. (2023) Gardner, J. P., Mather, J. C., Abbott, R., et al. 2023, PASP, 135, 068001
- Giménez-Arteaga et al. (2024) Giménez-Arteaga, C., Fujimoto, S., Valentino, F., et al. 2024, A&A, 686, A63
- GrudiÄ et al. (2019) GrudiÄ, M. Y., Hopkins, P. F., Quataert, E., & Murray, N. 2019, MNRAS, 483, 5548
- Harikane et al. (2018) Harikane, Y., Ouchi, M., Shibuya, T., et al. 2018, ApJ, 859, 84
- Harshan et al. (2024) Harshan, A., Tripodi, R., Martis, N. S., et al. 2024, ApJ, 977, L36
- Hashimoto et al. (2023) Hashimoto, T., Ălvarez-MĂĄrquez, J., Fudamoto, Y., et al. 2023, ApJ, 955, L2
- Hashimoto et al. (2019) Hashimoto, T., Inoue, A. K., Mawatari, K., et al. 2019, Publications of the Astronomical Society of Japan, 71
- Hashimoto et al. (2023) Hashimoto, T., Inoue, A. K., Sugahara, Y., et al. 2023, The Astrophysical Journal, 952, 48
- Hashimoto et al. (2018) Hashimoto, T., Laporte, N., Mawatari, K., et al. 2018, Nature, 557, 392
- Heintz et al. (2023) Heintz, K. E., Brammer, G. B., Giménez-Arteaga, C., et al. 2023, Nature Astronomy, 7, 1517
- Hu et al. (2025) Hu, W., Papovich, C., Shen, L., et al. 2025, arXiv e-prints, arXiv:2503.04032
- Hu et al. (2025) Hu, W., Papovich, C., Shen, L., et al. 2025, Circumgalactic Enrichment by a Multi-Galaxy Merger at Redshift 6.7
- Jones et al. (2024) Jones, G. C., Bowler, R., Bunker, A. J., et al. 2024, arXiv e-prints, arXiv:2412.15027
- Jones et al. (2023) Jones, O. C., Ălvarez-MĂĄrquez, J., Sloan, G. C., et al. 2023, MNRAS, 523, 2519
- Katz et al. (2024) Katz, H., Cameron, A. J., Saxena, A., et al. 2024, arXiv e-prints, arXiv:2408.03189
- Kennicutt & Evans (2012) Kennicutt, R. C. & Evans, N. J. 2012, ARA&A, 50, 531
- Kennicutt et al. (2009) Kennicutt, R. C., Hao, C.-N., Calzetti, D., et al. 2009, The Astrophysical Journal, 703, 1672
- Kohandel et al. (2024) Kohandel, M., Pallottini, A., Ferrara, A., et al. 2024, A&A, 685, A72
- Kokorev et al. (2024) Kokorev, V., Caputi, K. I., Greene, J. E., et al. 2024, ApJ, 968, 38
- Labiano et al. (2021) Labiano, A., Argyriou, I., Ălvarez-MĂĄrquez, J., et al. 2021, A&A, 656, A57
- Lam et al. (2019) Lam, D., Bouwens, R. J., Labbé, I., et al. 2019, A&A, 627, A164
- Langeroodi & Hjorth (2023) Langeroodi, D. & Hjorth, J. 2023, arXiv e-prints, arXiv:2307.06336
- Langeroodi & Hjorth (2024) Langeroodi, D. & Hjorth, J. 2024, arXiv e-prints, arXiv:2404.13045
- Law et al. (2023) Law, D. D., Morrison, J. E., Argyriou, I., et al. 2023, AJ, 166, 45
- Le FÚvre et al. (2020) Le FÚvre, O., Béthermin, M., Faisst, A., et al. 2020, A&A, 643, A1
- Leroy et al. (2021) Leroy, A. K., Schinnerer, E., Hughes, A., et al. 2021, ApJS, 257, 43
- Libralato et al. (2024) Libralato, M., Argyriou, I., Dicken, D., et al. 2024, Publications of the Astronomical Society of the Pacific, 136, 034502
- Maseda et al. (2020) Maseda, M. V., Bacon, R., Lam, D., et al. 2020, MNRAS, 493, 5120
- Matthee et al. (2017) Matthee, J., Sobral, D., Best, P., et al. 2017, MNRAS, 465, 3637
- Meurer et al. (1999) Meurer, G. R., Heckman, T. M., & Calzetti, D. 1999, ApJ, 521, 64
- Mihos & Hernquist (1996) Mihos, J. C. & Hernquist, L. 1996, ApJ, 464, 641
- Mowla et al. (2024) Mowla, L., Iyer, K., Asada, Y., et al. 2024, Nature, 636, 332
- Naidu et al. (2025) Naidu, R. P., Oesch, P. A., Brammer, G., et al. 2025, arXiv e-prints, arXiv:2505.11263
- Nakazato et al. (2024) Nakazato, Y., Ceverino, D., & Yoshida, N. 2024, ApJ, 975, 238
- Nardiello et al. (2022) Nardiello, D., Bedin, L. R., Burgasser, A., et al. 2022, MNRAS, 517, 484
- Navarro-Carrera et al. (2024) Navarro-Carrera, R., Rinaldi, P., Caputi, K. I., et al. 2024, arXiv e-prints, arXiv:2410.23249
- Ning et al. (2023) Ning, Y., Cai, Z., Jiang, L., et al. 2023, ApJ, 944, L1
- Noll et al. (2009) Noll, S., Burgarella, D., Giovannoli, E., et al. 2009, Astronomy & Astrophysics, 507, 1793
- Ăstlin et al. (2025) Ăstlin, G., PĂ©rez-GonzĂĄlez, P. G., Melinder, J., et al. 2025, A&A, 696, A57
- Peng et al. (2002) Peng, C. Y., Ho, L. C., Impey, C. D., & Rix, H.-W. 2002, AJ, 124, 266
- Pérez-Gonzålez et al. (2024a) Pérez-Gonzålez, P. G., Barro, G., Rieke, G. H., et al. 2024a, ApJ, 968, 4
- Pérez-Gonzålez et al. (2024b) Pérez-Gonzålez, P. G., Barro, G., Rieke, G. H., et al. 2024b, ApJ, 968, 4
- PĂ©rez-GonzĂĄlez et al. (2024c) PĂ©rez-GonzĂĄlez, P. G., DâEugenio, F., RodrĂguez del Pino, B., et al. 2024c, arXiv e-prints, arXiv:2405.03744
- PĂ©rez-GonzĂĄlez et al. (2025) PĂ©rez-GonzĂĄlez, P. G., Ăstlin, G., Costantin, L., et al. 2025, arXiv e-prints, arXiv:2503.15594
- Planck Collaboration et al. (2020) Planck Collaboration, Aghanim, N., Akrami, Y., et al. 2020, A&A, 641, A6
- Prieto-Lyon et al. (2023) Prieto-Lyon, G., Strait, V., Mason, C. A., et al. 2023, A&A, 672, A186
- Puskås et al. (2025) Puskås, D., Tacchella, S., Simmonds, C., et al. 2025, MNRAS, 540, 2146
- Reddy et al. (2018) Reddy, N. A., Shapley, A. E., Sanders, R. L., et al. 2018, ApJ, 869, 92
- Rieke et al. (2015) Rieke, G. H., Wright, G. S., Böker, T., et al. 2015, PASP, 127, 584
- Rieke et al. (2023) Rieke, M. J., Kelly, D. M., Misselt, K., et al. 2023, Publications of the Astronomical Society of the Pacific, 135, 028001
- Rinaldi et al. (2024a) Rinaldi, P., Bonaventura, N., Rieke, G. H., et al. 2024a, arXiv e-prints, arXiv:2411.14383
- Rinaldi et al. (2023) Rinaldi, P., Caputi, K. I., Costantin, L., et al. 2023, ApJ, 952, 143
- Rinaldi et al. (2024b) Rinaldi, P., Caputi, K. I., Iani, E., et al. 2024b, ApJ, 969, 12
- Rinaldi et al. (2025) Rinaldi, P., Navarro-Carrera, R., Caputi, K. I., et al. 2025, ApJ, 981, 161
- Rodriguez-Gomez et al. (2019) Rodriguez-Gomez, V., Snyder, G. F., Lotz, J. M., et al. 2019, MNRAS, 483, 4140
- Rowland et al. (2024) Rowland, L. E., Hodge, J., Bouwens, R., et al. 2024, MNRAS, 535, 2068
- Schaerer et al. (2020) Schaerer, D., Ginolfi, M., Béthermin, M., et al. 2020, A&A, 643, A3
- Speagle et al. (2014) Speagle, J. S., Steinhardt, C. L., Capak, P. L., & Silverman, J. D. 2014, ApJS, 214, 15
- Spilker et al. (2022) Spilker, J. S., Hayward, C. C., Marrone, D. P., et al. 2022, ApJ, 929, L3
- Spilker et al. (2023) Spilker, J. S., Phadke, K. A., Aravena, M., et al. 2023, Nature, 618, 708
- Stanway & Eldridge (2023) Stanway, E. R. & Eldridge, J. J. 2023, MNRAS, 522, 4430
- Stanway et al. (2016) Stanway, E. R., Eldridge, J. J., & Becker, G. D. 2016, MNRAS, 456, 485
- Stefanon et al. (2022) Stefanon, M., Bouwens, R. J., Illingworth, G. D., et al. 2022, ApJ, 935, 94
- Sugahara et al. (2025) Sugahara, Y., Ălvarez-MĂĄrquez, J., Hashimoto, T., et al. 2025, ApJ, 981, 135
- Sugahara et al. (2021) Sugahara, Y., Inoue, A. K., Hashimoto, T., et al. 2021, The Astrophysical Journal, 923, 5
- Sun et al. (2023) Sun, F., Egami, E., Pirzkal, N., et al. 2023, ApJ, 953, 53
- Toomre & Toomre (1972) Toomre, A. & Toomre, J. 1972, ApJ, 178, 623
- Wells et al. (2015) Wells, M., Pel, J. W., Glasse, A., et al. 2015, PASP, 127, 646
- Wright et al. (2023) Wright, G. S., Rieke, G. H., Glasse, A., et al. 2023, PASP, 135, 048003
- Wright et al. (2015) Wright, G. S., Wright, D., Goodson, G. B., et al. 2015, PASP, 127, 595